OpenStudy (anonymous):

If $360 is invested at an interest rate of 4% per year and is compounded quarterly, how much will the investment be worth in 18 years?

OpenStudy (anonymous):

\[A=P\left(1+{r\over 100}\right)^n\]

OpenStudy (anonymous):

P=360 r=4 n=18 find A

OpenStudy (anonymous):

wait.. quatrterly

OpenStudy (anonymous):

so, \[A=P\left(1+{r\over 100k}\right)^{kn}\] k = 4 ( 4 quarters per year)

OpenStudy (anonymous):

kapeesh?

OpenStudy (anonymous):

\[360\times \left(1+\frac{.04}{4}\right)^{4\times 18}\] and a calculator

OpenStudy (anonymous):

736.96..Thanks guys!! :D

OpenStudy (anonymous):

yw